OpenStudy (anonymous):

Find 3 consecutive odd numbers where the product of the smaller two numbers is 34 less than the square of the largest number. @alli14344

OpenStudy (anonymous):

3,5,7

OpenStudy (anonymous):

x > y > z x - 2 = y y - 2 = z yz = x^2 - 34 x - 2 = y x = y + 2 yz = x^2 - 34 y(y - 2) = (y + 2)^2 - 34 y^2 - 2y = (y + 2)(y + 2) - 34 y^2 - 2y = y^2 + 2y + 2y + 4 - 34 y^2 - y^2 - 2y - 4y = 4 - 34 -6y = -30 y = -30/-6 y = 5 x - 2 = y x - 2 = 5 x = 5 + 2 x = 7 y - 2 = z 5 - 2 = z z = 3 ∴ x = 7 , y = 5 , z = 3 or you could do it this way: suppose the smallest odd number was x the second number would be x + 2, and the third number would be x + 4 so the product of the two smaller numbers (x, x + 2) is 34 less than the square of the largest number (x + 4) x(x+2) = (x+4)^2 - 34 x^2 + 2x = x^2 + 8x + 16 - 34 6x = 18 x = 3 so the three number are 3, 5, and 7